OpenStudy (anonymous):

P(x) = x^2 + x – 2. Find P(2). A. 0 B. 2 C. 4 D. 7 is this B??

OpenStudy (bradely):

P(x) = x^2 + x – 2 P(2)=2^2+2-2 =4+2-2 =4 Option C is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

ok thanks